OpenStudy (grace111):

Please help me!!!! Will give medal and fan for best answer. !! What is 30% of £200? What is 1/4 of 20?

jhonyy9 (jhonyy9):

try calculi it using the rule of 3 simple so : 100% ----- 200 30% ---- x ---------------------- than cross multiplie and calculi x 100x = 30*200 so than x = 30*200 /100 = ?

OpenStudy (grace111):

60?

jhonyy9 (jhonyy9):

i think yes

OpenStudy (grace111):

oh thank you

jhonyy9 (jhonyy9):

my pleasure good luck bye

OpenStudy (grace111):

bye